OpenStudy (anonymous):

A cylindrical vase has a diameter of 4 inches. At the bottom of the vase, there are 6 marbles, each of diameter 2 inches. The vase is filled with water up to a height of 8 inches. What is the volume of water in the vase? 26π in3 16π in3 24π in3 32π in3

OpenStudy (anonymous):

c

OpenStudy (graciebugg):

I've seen this question before and I'm pretty sure it's D

OpenStudy (anonymous):

no it 24pi

OpenStudy (graciebugg):

Substract the volume of the marbles 32π−8π=24π

OpenStudy (anonymous):

Volume of water = Volume of water 10 inch high - Volume of 6 marbles = π(3 in)^2 * (10 in) - 6(4/3)(π)(1 in)^3 = 90π in^3 - 8π in^3 = 82π in^3 and this is how you get it

OpenStudy (anonymous):

So it's 24?

OpenStudy (graciebugg):

Yeah I see where I missed a letter Its 24

OpenStudy (anonymous):

whut grade r u in lol jw @codyisbeast

OpenStudy (anonymous):

bc i am in 9th